### v1

# Dynamical complex networks for understanding motility induced phase separation in Active Brownian Particles

Abstract
We investigate the dynamics of active Brownian particles (ABP) through a temporal complex network framework, focusing on degree probability distribution, average path length, and clusterization across Motility-Induced Phase Separation (MIPS) regions. In the single MIPS phase, particle encounters resemble a random graph. Transitioning toward the MIPS frontier introduces a non-power-law distribution, while the phase-separated region exhibits a double power-law distribution. Numerical analysis indicates a shift from random to power-law complex networks based on the system’s MIPS-phase position. In summary, our numerical and theoretical analysis reveals the phase transition of the system between two states and how this transition manifests itself in the complex network, showing an evident change between a random network and a network that tends to have scale-free behavior.
1 Introduction
Active Brownian Particles (ABP) serve as a valuable framework for investigating the non-equilibrium dynamics and statistical properties of self-propelled agents, whether they are living organisms or artificial entities romanczuk2012active ; bechinger2016active . In contrast to passive Brownian particles studied in random diffusive systems, ABP offers a new dimension of complexity due to the inherent particle self-propulsion mechanisms. In particular, these active particles undergo a distinctive phase transition known as Motility-Induced Phase Separation (MIPS) fily2012athermal ; redner2013structure ; stenhammar2014phase ; theers2018clustering ; van2019interrupted . This phase transition is related to the spontaneous separation between dilute (free-particles) and dense phases (cluster formation) stenhammar2014phase ; de2023sequential ; de2021active . This phase separation is a fascinating emergent phenomenon where particles collapse into clusters although natural repulsion forces are present. The exploration of ABP not only deepens our understanding of self-organization and emergent phenomena but also opens up promising avenues for harnessing their unique properties in various applications and technological advancements gompper20202020 and building a bridge between physical models and biological systems.
The complexity of ABP dynamics and its statistical properties offer novel opportunities for deeper complex network analysis to better understand the role of particle connections during the dynamics. The particle interaction or contact dynamic between particles within this system is still a significant topic of interest, where the effects of particle shape theers2018clustering , inertia caprini2022role , hydrodynamic interactions zottl2023modeling ; theers2018clustering ; tan2022odd ; petroff2015fast and quorum sensing ridgway2023motility among others has been studied numerically and analytically in the context of motility-induced clustering and motility-induced separation. In recent years agent-based-models such as ABP have been used to construct from particle direct physical contact the contagion dynamics on the top of compartmental ecological models norambuena2020understanding ; de2023sequential ; forgacs2022using , or from more complicated pathways where temporal networks has been used to track burstiness or clusterization in the system richardson2015beyond ; gorochowski2017behaviour ; zhong2023burstiness ; bhaskar2021topological . In recent work, McDermott et al. unveil the time evolution hidden in the clustering effect during MIPS using machine learning mcdermott2023characterizing , where they demonstrate different regimes close the phase separation.
In the last years, studies conducted by complex networks have been applied to many natural systems such as social interactions newman2001 ; Kertesz ; gonzalez2008understanding , seismology abe2006 ; telesca2012 ; pasten2016 ; pasten2018 , space plasmas suyal2014visibility ; mohammadi2021complex , or biological interactionsthiery ; barabasi2011 ; scabini ; riley2007large ; stockmaier2021infectious , among others. The formalism of complex networks can show a non-trivial behavior in systems due to their geometrical and topological foundations newman2003 ; albert2002statistical . Throughout a complex network analysis, we can characterize and describe a physical system based on its topological properties. Thus, we can classify a complex network based on its structural organization of connections between nodes such as a random network erdos1960 , a scale-free network barabasi1999 , or a small-world network watts1998 , for example. Due to the random encounters between ABP, one can construct a temporally averaged network that allows us to analyze the dynamics in terms of the topology of such connections. This new approach may motivate new questions or characterize the phase transition (MIPS) using the language of complex networks.
In this work, we carry out a new manner of analyzing ABP using complex networks: we build a time-averaged network with the time evolution of a system of ABP, and we follow the changes in the topological features of the complex network through the degree distribution ($P(k)$), the clustering coefficient ($\bar{C}$) and the average path length ($l_{G}$). Our approach suggests that a random graph can characterize the dilute phase while a double power law distribution describes the MIPS region.
2 Active Brownian Particles (ABP) and motility-phase-separation
We consider $N$ self-propelled disk-like particles with position vectors $\mathbf{r}_{i}(t)=[x_{i}(t),y_{i}(t)]$ (center of the disk) and orientation vectors $\hat{\mathbf{u}}_{i}(t)=[\cos\theta_{i}(t),\sin\theta_{i}(t)]$ that obey overdamped Langevin dynamics in a flat 2D surface,
$\dot{\mathbf{r}}_{i}$ $=$ $u_{0}\hat{\mathbf{u}}_{i}(t)+\bm{\xi}_{i,T}(t)/\gamma_{T}-(1/\gamma_{T})\nabla_{\mathbf{r}_{i}}U,$ (1)
$\dot{\hat{\mathbf{u}}}_{i}$ $=$ $(1/\gamma_{R})\bm{\xi}_{i,R}(t)\times\hat{\mathbf{u}}_{i}$ (2)
where $\gamma_{T},\gamma_{R}$ are the Stokes friction coefficients associated with translation (T) and rotation (R) in the $x-y$ plane, respectively, $u_{0}$ represents the self-propelled velocity with constant magnitude see Figure 1. Also, in Eqs.(1) and (2), $\bm{\xi}_{T}(t)$ and $\bm{\xi}_{R}(t)$ are, respectively, the translational and rotational, Gaussian, white noises with zero mean, $\langle\bm{\xi}_{T}(t)\rangle=\langle\bm{\xi}_{R}(t)\rangle=\mathbf{0}$, and time correlation $\langle\bm{\xi}_{T}(t_{1})\bm{\xi}_{T}(t_{2})\rangle=D_{T}\delta(t_{1}-t_{2})$ and $\langle\bm{\xi}_{R}(t_{1})\bm{\xi}_{R}(t_{2})\rangle=D_{R}\delta(t_{1}-t_{2})$, where we can define $D_{T}$ and $D_{R}$ as the translation and rotation diffusion parameters. Finally, in Eq.(1), $U$ corresponds to the summation over steric interactions between particles at time $t$,
$U=\sum_{i\neq j}U_{\text{WCA}}(\mathbf{r}_{i},\mathbf{r}_{j})$ (3)
with $\mathbf{r}_{j}$ is the position of the $j-$th particle, and $U_{\text{WCA}}$ is the Weeks-Chandler-Andersen (WCA) pair potential
$U_{\text{WCA}}(\mathbf{r}_{i},\mathbf{r}_{j})=\left\{\begin{array}[]{cc}\displaystyle{4\varepsilon\left[\left(\frac{\sigma}{r_{ij}}\right)^{12}-\left(\frac{\sigma}{r_{ij}}\right)^{6}\right]}&r_{ij}\leq C_{r}\\ 0&\mbox{otherwise}\end{array}\right.$ (4)
with $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ is the distance between particles, $\sigma$ the particle diameter, $C_{r}=2^{1/6}\sigma\approx 1.1225\sigma$ is defined here as the contact radius, and $\varepsilon$ is the energy interaction strength between particles. We set $\varepsilon=3$ during our simulations unless we stated otherwise.
Figure 1: Time evolution for a system of active Brownian particles exhibiting Motility-induced phase separation: In the top-left, a single ABP is shown with orientation vector $\hat{u}_{i}$ and radius $R$. The time evolution is shown at three different times, showing a transition from a single phase in $t=0.02t_{r}$ and two coexisting phases in $t=5.85t_{r}$, at $\text{Pe}=140$ and $\phi=0.5$, with $L=100$ and $t_{r}=1/D_{R}$.
In this study, we conducted numerical simulations involving a collection of ABP moving within a square of area $L\times L$. Here, it is important to introduce the two dimensionless numbers to characterize different regimes. First, we introduce the packing fraction as $\phi=N\pi(\sigma/2)^{2}/L^{2}$, which relates the ratio between the area covered by the $N$ Active Brownian Particles with radius $\sigma/2$ with respect to the total available area. When the packing fraction increases, the high density allows clusters to be formed. In addition, these ABP systems exhibit a characteristic persistent motion reminiscent of a run-and-tumble state observed in self-phoretic systems and bacterial suspensions cates2015motility ; buttinoni2013dynamical ; petroff2015fast ; van2019interrupted . The persistence time is defined as $t_{p}=\sigma/u_{0}$, while the memory kernel associated with tumble states is characterized by $t_{r}=1/D_{R}$. Since ABP can explore the available space through diffusion, it is natural to define $t_{r}$ as the relevant time scale of the problem. Second, we introduce the Péclet number $\text{Pe}=t_{r}/t_{p}=u_{0}t_{r}/\sigma$, which relates the ratio between diffusion and persistent times in transport phenomena. When the Péclet number is large, we have that ABP are persistent compared to diffusion. The latter is relevant for cluster formation since the minimal cluster for disk-shaped particles is formed with three persistent disks forming an equilateral triangle.
The interplay between steric interactions causing the blocking of persistent motion and an increased probability of collision due to elevated concentrations $\phi$ and particle activities $u_{0}$ leads to a fascinating phenomenon known as motility-induced phase separation (MIPS) theers2018clustering ; stenhammar2014phase . The system starts with a homogeneous particle distribution and orientations at $t=0$ exhibiting a single-phase behavior (or fluid-like state su2023motility ), which physically describes ABP randomly moving without particle agglomeration. However, if the packing fraction and Péclet number are larger than some critical values, the system evolves into a two-phase state represented by a solid and gas phase, as illustrated in Figure 1. In this phase, particles agglomerate, and cluster formation becomes an emergent phenomenon.
An analytical alternative to represent this phase separation is using the separation kinetic function $f_{c}$, which reproduces the phase separation map in terms of the packing fraction, the Péclet number, and the average total number of particles lost per escape event $\kappa$ redner2013structure
$f_{c}=\frac{4\phi\text{Pe}-3\pi^{2}\kappa}{4\phi\text{Pe}-6\sqrt{3}\pi\kappa\phi}$ (5)
In Fig. 2, we analytically reproduced the phase diagram in terms of the packing fraction and the Péclet number using $\kappa=4.7$ as a fit parameter redner2013structure . The pairs $(\text{Pe},\phi)$ studied in this work are represented by the white dots of Fig. 2. For $f_{c}=0$, we can say that the system is in the single-phase, while for $f_{c}\neq 0$, the system undergoes a phase transition to the two-phase state. The system is considered steady when the phase separation shows a bimodal probability distribution for different local packing fractions redner2013structure ; stenhammar2014phase . Despite extensive studies on phase separation in various contexts, machine learning has recently investigated the time evolution during phase separation, revealing different transition regimes mcdermott2023characterizing . In this work, we aim to study the time evolution by generating a temporal averaged network during the phase transition.
Figure 2: An analytical representation for the phase diagram in the $\text{Pe}-\phi$ plane plotted from the analytical model using Eq. 5, with $\kappa=4.7$. The white dots represent the pairs $(\text{Pe},\phi)$ parameters studied across this work.
2.1 Simulation methods and parameters
We employed standard Brownian dynamics methods to numerically solve the governing Eqs.(1) and (2) for $N\in[955,1274,1592,1910,2229]$ particles within a simulation box with a length of $L=100$ and periodic boundary conditions. We ran $20$ ensembles for each set of parameters, initiating simulations with random initial particle positions and orientations.
The relevant length scale in the system is the particle radius $R$, serving as the unit of length. Additionally, we adopted $t_{r}$ as the unit of time. The equations of motion, Eqs.(1) and (2), were integrated with a time step of $\Delta t=3\times 10^{-5}t_{r}$, and simulation times were set to $T=12t_{r}$. Two dimensionless values parameterize the system: the packing fraction $\phi$ and the Péclet number, which are varied in the intervals $\phi\in[0.3-0.7]$ and $\text{Pe}\in[80,140]$, respectively (see Fig. 2). We maintain a constant rotational diffusion coefficient $D_{R}=0.15$, translation diffusion coefficient $D_{T}=0.2$, and diameter $\sigma=2R$ with $R=1$.
3 Complex Networks for ABP
A network is defined as a group of nodes connected between them by edges. The connections between nodes could be directed or undirected. In this case, the nodes are determined by each ABP, and the interaction between particles defines the connections: two nodes (particles) are connected when their centers are inside the interaction region defined by the contact radius $C_{r}$ (see WCA potential in Eq. (4)). We build a time network by following the dynamics of the particle connections in a specific time window.
To have enough statistics, we define a window with sixty data points, and we count the connections between nodes each time; if two nodes are connected only once in a one-time step or if two nodes are connected along the sixty steps, the value of the connections in the adjacency matrix is one in that time window. We move the window to the next time window’s center and repeat the same procedure, as shown schematically in Fig. 3.
Figure 3: From left to right: A scheme for particle encounters when a pair interaction occurs. The blurred area represents the interaction region with radius $C_{r}$. Particles (nodes) generate a link if their centers enter each other’s interaction region, as shown in the middle panel. A representation of the time evolution for the network formation, with time windows of length $W=0.45t_{r}$ and overlap $t_{0}=0.225t_{r}$ between two consecutive windows.
3.1 Adjacency matrix
In order to analyze the system, we first create the adjacency matrix $A$, which is used to represent the network. Using the distance $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ between the centers of two particles, we can establish the value of $A_{ij}$ which represents the connections between these particles. If $r_{ij}\leq C_{r}$, we set $A_{ij}=A_{ji}=1$; otherwise, $A_{ij}=0$. If $A_{ij}=A_{ji}=1$ it means that two nodes (particles) are connected between them, and the complex network built is undirected. Within time windows of length $W=0.45t_{r}$ and using an overlapping time $t_{0}=0.225t_{r}$, a link is established only upon the first interaction between two particles. The latter means that each network accounts for unique particle encounters within the time window. From the adjacency matrix, we can compute all the necessary properties of the network. We will focus on calculating the degree distribution, network average clustering coefficient, and average path length.
3.2 Degree distribution
The degree $deg(p_{i})$ of each particle $p_{i}$ is the number of connections of a node, and it can be computed by summing over the corresponding column or row of the adjacency matrix:
$deg(p_{i})=\sum_{j=1}^{N}A_{ij}=\sum_{j=1}^{N}A_{ji},$ (6)
and the degree distribution is defined by
$P(k)=\frac{n_{k}}{N},$ (7)
which is directly obtained from simulations. Here $n_{k}$ is the total number of particles with degree $k$, and $N$ is the number of particles in the system. Depending on the topology of the complex network, the degree distribution could be binomial (random graphs), power-law (scale-free), among others. In this work, we shall discuss the role of the Gaussian distribution and double power law models in the context of MIPS.
3.3 Network average clustering coefficient
The clustering coefficient $C_{i}$ of each particle is a measure of the tendency to form clusters of this node, and it is defined by
$C_{i}=\frac{1}{k_{i}(k_{i}-1)}\sum_{j,k}A_{ij}A_{jk}A_{ki},$ (8)
where $A$ is the adjacency matrix, and $k_{i}$ is the degree of the particle. From this, we can calculate the network average clustering coefficient $\bar{C}$, which is the trend of the network to form clusters, and it is defined by
$\bar{C}=\frac{1}{N}\sum_{i=1}^{N}C_{i},\quad\quad 0\leq\bar{C}\leq 1.$ (9)
3.4 Average path length
The average path length is the rate between the shortest number of steps between two nodes and the shortest paths between all the pairs of nodes, and it is defined by Newman_2000 ,
$l_{G}=\frac{1}{N(N-1)}\sum_{i\neq j}d(v_{i},v_{j}).$ (10)
Here, $d(v_{i},v_{j})$ is the shortest distance between two vertices $v_{i}$ and $v_{j}$, where $v_{i,j}$ are nodes of the graph.
4 Results and Discussion
We studied three different regimes of the ABP according to the phase diagram of Fig. 2: (i) one-phase (violet), (ii) transient phase (blue-green), and (iii) two-phase (orange-red). As discussed in the following sections, these three regions will exhibit different topological properties, which are not reported in the literature to the best of our knowledge.
4.1 One-phase region: Random Graph
In the one-phase region (violet in Fig. 2), we found that the random motion of ABP can be characterized by a Gaussian function for the degree distribution
$P^{\rm G}(k)={\frac{1}{\sigma_{\rm st}\sqrt{2\pi}}}\mbox{exp}\left[-\frac{1}{2}\left(\frac{k-\mu}{\sigma_{\rm st}}\right)^{2}\right],$ (11)
where $\sigma_{\rm st}$ is the standard deviation, and $\mu$ is the average number of unique encounters of particles in a time window $W$. We have a dilute system for packing fraction numbers $\phi<0.4$, which allows us to analyze the average behavior of the ABP system using the kinetic theory of gases. As explained in Ref. norambuena2020understanding , in an active media with $N$ moving particles in a dilute system, the mean free path is $\lambda=\sqrt{2}u_{0}\tau_{c}$, where $\tau_{c}$ is the mean time between encounters. During a time interval $\tau_{c}$, we can establish the probability constraint $p_{c}=A_{\rm sw}/L^{2}\approx 1$, where $A_{\rm sw}=N(2\lambda a+\pi a^{2})$ is the total area swept area of the $N$ particles, and $L^{2}$ is the available area. From this argument, it follows that $\lambda=(\pi/4)\phi^{-1}(1-\phi)\sigma=\sqrt{2}u_{0}\tau_{c}$. Then, the average number of encounters can be estimated as $\mu=W/\tau_{c}$, since, on average (by definition), particles have one encounter in the time interval $\tau_{c}$. We found:
$\mu(\phi,\text{Pe})={4\sqrt{2}\over\pi}\left({\phi\over 1-\phi}\right)WD_{R}\text{Pe},$ (12)
Figure 4: Degree distribution $P(k)$ for systems in the one-phase region at $\text{Pe}=80$ (circles) and $\text{Pe}=140$ (diamonds) at time $t=0.45t_{r}$ (top) and at $t=11.0t_{r}$ (bottom). The distributions show a Gaussian tendency; each system exhibits distinct mean and standard deviation values. Note that all systems maintain a Gaussian distribution, and only the system at $\phi=0.4$ shows a change over time. Both insets (top and bottom) show the $\log-\log$ plot of the systems at the respective time.
where $W=0.45t_{r}$ is fixed for $D_{R}=0.15$. In Fig. 4, we illustrate the evolution of the degree distribution $P(k)$ measured from Eqs. 6 and 7 for a system at two time lapses, $t=0.45t_{r}$ and $t=11.0t_{r}$. Two Péclet numbers are considered: $\text{Pe}=80$ represented by circles and $\text{Pe}=140$ represented by diamonds. Different packing fractions are depicted with orange and blue colors, corresponding to $\phi=0.3$ and $\phi=0.4$, respectively. Most importantly, using our model given in Eq. 12, we obtain $\mu(\phi=0.3,\text{Pe}=80)\approx 10.3$ and $\mu(\phi=0.3,\text{Pe}=140)\approx 16.2$, whose numerical values are close to those obtained by simulations ($10.51$ and $16.75$), see Fig. 5. This simple model can give good estimations of the central value of the Gaussian distribution $P^{\rm G}(k)$ given in equation 11, which is valid for the one-phase region.
Figure 5: Standard deviation $\sigma_{st}$ (top) and mean $\mu$ (bottom) values over time of Gaussian fits for the degree distribution of systems in the one-phase region (Fig. 4). Both $\sigma_{st}$ and $\mu$ remain virtually constant over time and only the mean for the system at $\phi=0.4$ shows an appreciable decrease.
4.2 Transient Region: Hybrid phase
We define the transient region as the intermediate phase between the one-phase and two-phase regions (blue-green in Fig. 2). Since it is an intermediate phase, it is expected to have a degree of distribution with a hybrid behavior. Physically and analyzing our numerical simulations, we note that in this regime, small clusters composed of three disks (triangular arrangement) appear and eventually break because of the persistent ABP motion. In Fig. 6, we observe the degree distribution, average path length ($l_{G}$), and average clustering coefficient ($\bar{C}$) at times $t=0.45t_{r}$ and $t=11t_{r}$. Interestingly, the initially random graph (Gaussian distribution) evolves to a hybrid distribution with two components (Gaussian and power law). This hybrid behavior tells us that the system is trying to evolve from a random to a scale-free graph, but the values for the packing fraction and Péclet are not large enough. Moreover, at $t=11t_{r}$, the system does not reach the steady state, which is not the case for the one-phase and two-phase regions. The latter tells us that this hybrid phase is more complex to analyze regarding its geometrical and topological properties since the system remains out of equilibrium.
Figure 6: Top: Degree distribution for the case at $\text{Pe},\phi=80,0.4$ at the initial time $t=0.45t_{r}$ (right) and time $t=11.0t_{r}$ (left). The inset shows the same data plotted in $\log{}-\log{}$. Bottom: Average path length (left ) and average clustering coefficient (left) over time of the system. Both show an increasing behavior over time until $\approx 8t_{r}$.
4.3 Two-phase region: Phase separation
At more significant packing fractions and Péclet numbers, i.e., $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$, the system enters the inner region of Fig. 2, where MIPS is manifested. From the point of view of the ABP, as time evolves, the initially random distribution (liquid-like state) evolves to a state where a gas-like and a solid-like phase cohabit in the system (green-red regions in Fig. 2). In terms of the complex network, the degree distribution of the system evolves from a Gaussian distribution at short times ($t=0.45t_{r}$) towards a double power law distribution for larger times ($t\geq 11t_{r}$); see Fig. 7(a)(b). Notably, when the degree distribution is plotted in the log-log scale, two slopes adequately describe the distribution $P(k)$, see Fig. 7(c)(d). This double power law distribution has the form
$P(k>K)\sim\left\{\begin{array}[]{cr}k^{-\gamma_{1}},&k<k_{c}\\ k^{-\gamma_{2}},&k\geq k_{c}\end{array}\right.,$ (13)
where $K=6$, $k_{c}\in[13-17]$ is the critical degree and $(\gamma_{1},\gamma_{2})$ are the critical exponents. The value $K=6$ is connected with the 2D disk packing problem, where each disk has exactly six nearest neighbors in an ideal cluster and corresponds to the peak of $P(k)$, see Fig. 7(b). Moreover, the critical values for $k_{c}$ depend on the clustering coefficient and Péclet number. The first slope $\gamma_{1}$ (red dashed line of Fig. 7) changes from values close to $1.5$ at $\phi=0.5$ towards values close to $3.7$ at $\phi=0.7$. The second slope $\gamma_{2}$ (black dashed lines) varies from values close to $12$ to $17$. In summary, we obtain $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$, and therefore $\gamma_{1}<\gamma_{2}$ in our ABP model. In Table 1, we compare our results with different systems that exhibit a double power law distribution.
Physically, the $\gamma_{1}$ exponent is related to the solid cluster appearance where the degree number has a peak at $k=K$ ($\mbox{log}_{10}(6)\approx 0.78$). To better understand this, we plot the time evolution in Fig. 8, where it is clear that the slope value and, therefore, the cluster solid-like region is determined strongly by the system packing fraction $\phi$. On the contrary, the $\gamma_{2}$ exponent is related to less probable encounters related to the gas region, and we will not delve into that aspect.
The temporal behavior of both curves is shown in Fig. 8 ($\gamma_{1}$) and ($\gamma_{2}$). We note that $\gamma_{1}$ is more stable over time, and the packing fraction dominates its steady value in some inverse relation. On the contrary, $\gamma_{2}$ shows more oscillations, which can be explained because the clustering formation is unstable for the values of $(\phi,\mbox{Pe})$.
Table 1: Comparison between critical exponents for the double power law distribution in different models. Here, we are using the acronyms CAN = Chinese Airline Network, ABP = Active Brownian Particles, and MIPS = Motolity-Induced Phase Separation.
System Network $\gamma_{1}$ $\gamma_{2}$
or Data (nodes, edges) approx. approx.
CAN Data Li2004 (airports,flights) $0.41-0.45$ $4.16-4.54$
US Flight Network Data Li-Ping:1393 (airports,flights) $0.56-0.57$ $3.27-3.41$
Evolution of CAN Han_2011 (airports,flights) $0.46-0.52$ $2.1-2.7$
Evolution of CAN ZHANG20103922 (airports,flights) $0.49$ $2.63$
Language model Dorogovtsev2001 (words, connected words) $3/2$ $2$
Homer’s Odyssey Data PINTO20144019 (words, connected words) $0.76$ $1.19$
ABP (our work) (particles, interacting particles) $1.5-3.7$ $12-17$
Figure 7: Degree distribution for phase-separated systems. (a) $\log{}-\log{}$ plot shows a clear Gaussian distributions at time $t=0.45t_{r}$. (b) degree distribution at $t=11.0t_{r}$ shows an evolution in the distribution where the degree $k=6$ for each curve is the one with the highest probability. Figures (c) and (d) show two distributions of the figure (b) zoomed in at $\text{Pe},\phi=(80,0.7)$ and $\text{Pe},\phi=(140,0.5)$ respectively, where the red and black dashed lines are the best linear fit for the first and second power law decay respectively for each distribution.
Figure 8: Slopes of the power law decay in the degree distribution for phase-separated systems. The initial data point occurs when the degree $k=6$ has the highest probability, indicating distinct curves for each system. Consequently, we observe different transition times. Notably, around $t\approx 8t_{r}$, the value of $\gamma_{1}$ (top) stabilizes consistently across all systems. In contrast, the values for the second slope, $\gamma_{2}$ (bottom), appear noisy and do not seem to stabilize.
4.3.1 Role of the average clustering coefficient
We can characterize this phase separation using the network average clustering coefficient $\bar{C}$ given in Eq. 9. In Fig. 9, we plot $\bar{C}$ as a function of time for different values of the packing fraction and Péclet number for the frontier and inner regions of the phase diagram illustrated in Fig. 2. To better understand the observed values of $\bar{C}$, we can elaborate a simple theoretical analysis. The clustering coefficient can be computed as $C=\lambda_{G}(v)/\tau_{G}(v)$, where $\lambda_{G}(v)$ is the number of triangles with one vertex on $v$ and $\tau_{G}(v)$ is the number of triples with one vertex on $v$ ($v$ is a vertex of the graph $G$). For a simple hexagonal cluster composed of seven particles (see inset of Fig. 10 (left)), we can calculate the clustering coefficient for one particle inside the cluster ($C_{i}$) and one particle in the border ($C_{b}$). Using this reasoning, we obtain $C_{i}=6/15=0.4$ and $C_{b}=2/3\approx 0.6667$. Let be $\bar{C}_{hex,N_{r}}$ the average clustering coefficient for a hexagonal arrangement with $N_{r}$ concentric hexagonal rings. For $N_{r}$ hexagonal rings, we have $6N_{r}$ particles in the border and $1+3N_{r}(N_{r}+1)$ particles in total. Therefore, the fraction of particles in the border and inside the cluster are $f_{b}(N_{r})=6N_{r}/(1+3N_{r}(N_{r}+1))$ and $f_{i}=(1+3N_{r}(N_{r}-1))/(1+3N_{r}(N_{r}+1))$, respectively. Thus, we can define the average clustering coefficient as,
$\bar{C}_{hex,N_{r}}=f_{b}(N_{r})C_{b}+f_{i}(N_{r})C_{i}={2\over 15}\left[5-2f_{i}(N_{r})\right],$ (14)
where we have used $f_{b}+f_{i}=1$. For instance, for $N_{r}=1$ (seven particles) we get $\bar{C}_{hex,7}=(1/7)C_{m}+(6/7)C_{b}\approx 0.6286$ (first value of Fig. 10 (left)). However, as one increases the fractions of particles inside the cluster ($f_{i}\rightarrow 1$), we get $\bar{C}_{hex,N_{r}}\rightarrow C_{i}=0.4$ (asymptotic value of Fig. 10 (left)). Also, it is important to mention that three disks forming a minimal cluster have an average cluster coefficient $\bar{C}=1$ (maximum value).
In Fig. 9, we note that in the frontier region of the MIPS phase diagram ($\phi=\{0.3,0.4\}$ and Pe =$=\{80,140\}$), where small triangular clusters randomly appear, we observe $0.2\lesssim\bar{C}\lesssim 0.35$, and we associate these values for very low probability events (probabilities $\sim(0.2-0.35)$) of forming small triangular clusters. However, as we enter the inner region of the MIPS phase diagram ($\phi=\{0.5,0.6,0.7\}$ and Pe =$=\{80.140\}$), finite hexagonal clusters are dominant, explaining the observed values $0.35\lesssim\bar{C}\lesssim 0.42$ at large times oscillates around $0.4$. A steady state value $\bar{C}\approx 0.35$ can be explained by the fact that a perfect hexagonal cluster with a clustering coefficient of 0.4 can be penalized by the fraction of particles with respect to the total number of particles. A value $\bar{C}\approx 0.42$ can be understood as the reference value 0.4 with a small contribution of small triangular clusters.
Figure 9: Top: Network average clustering coefficient for systems in the frontier region (top) and the inner region (bottom). The values for $\bar{C}$ for one-phased systems(top) at $\phi=0.3$ have virtually no change over time and slowly increase at $\phi=0.4$. In contrast, $\bar{C}$ for the two-phased systems increases rapidly at first and then stabilizes around $t=8t_{r}$, showing a tendency to $\bar{C}=0.4$.
Figure 10: Left: Network average clustering coefficient of identical circles in a hexagonal packing arrangement, using $C_{r}=2^{1/6}\sigma$ ( black circles). $\bar{C}$ approaches asymptotically to $0.4$ (dashed red line), which is the value of the clustering coefficient of the center particle in the seven-particle arrangement (inset). Right: Average path length for the same hexagonal packing arrangement as a function of the packing fraction $\phi$ (black circles). We see that $l_{G}$ grows as $30\sqrt{\phi}$ (red dashed line) and could represent an upper bound with respect to $\phi$ for clustered ABP systems.
4.3.2 Role of the average path length
The average path length defined in Eq. 10 was measured for both regions, the frontier and inner region. In Fig. 11, we observe the time evolution for the network average path length in the case of a system with a single phase characterized by a Gaussian degree distribution $l_{G}$ has a constant value for different Pe numbers, meanwhile on the transient region for $\phi=0.4$ and Pe=$80$ the value increases slowly not reaching a constant value see blue dots in Figure 11.
In Fig. 10, we can observe a particular scaling for the average path length. In the ideal case (perfect hexagonal clusters with $N$ particles), we note that $l_{G}=30\phi^{0.5}$ matches our numerical calculations shown in Figure 10. However, our simulations shown in Fig. 11 suggest a different scaling for the steady state value of $l_{G}$, where $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$. We propose the following phenomenological expression
$l_{G}(\mbox{Pe},\phi)=\zeta(\mbox{Pe})\times\phi^{1.3},$ (15)
where $\zeta(\mbox{Pe}=80)=21$ and $\zeta(\mbox{Pe}=140)=8.5$ are values obtained to fit our data. Using this simple model, we found the following estimations $l_{G}(\mbox{Pe}=80,\phi=0.5)\approx 8.53$, $l_{G}(\mbox{Pe}=140,\phi=0.5)\approx 7.51$, $l_{G}(\mbox{Pe}=80,\phi=0.6)\approx 10.81$, $l_{G}(\mbox{Pe}=140,\phi=0.5)\approx 9.53$, $l_{G}(\mbox{Pe}=80,\phi=0.7)\approx 13.21$, $l_{G}(\mbox{Pe}=140,\phi=0.7)\approx 11.64$, which are in very good agreement with our numerical calculations. For the inner region, we observe a time evolution in the $l_{G}$ value where the average path length increases with particle concentration $\phi$ and reaches a constant value after $t\geq 8t_{r}$ which corresponds to the cluster growth characteristic time reported by Stenhammar et al.stenhammar2014phase .
Figure 11: Average path length for systems in the frontier region (top) and in the inner region (bottom). The values for $l_{G}$ for single-phase systems(top) at $\phi=0.3$ have virtually no change over time and at $\phi=0.4$ increases slowly. In contrast, $l_{G}$ for the separated phase increases rapidly at first and then stabilizes around $t=8t_{r}$.
5 Conclusions
In this work, we have studied the dynamics of ABP using complex networks. We build a time-averaged network and concentrate on the evolution of the topology of the complex network across the MIPS region. We follow the time evolution of the degree distribution, the clustering coefficient, and the average path length, finding different structures of the complex network along the time evolution of the ABP. In the one-phase region, we developed a numerical and theoretical analysis to explain the Gaussian distribution for random graphs. Using a mean free path analysis, we found an analytical expression for the average number of unique encounters of particles in a time window $W$, showing the role of the packing fraction and Péclet number. We reported a transient region, where the degree distribution is hybrid in the sense that it is represented by two components, Gaussian and power laws, namely. Also, in the hybrid region, the average path length and average clustering coefficient do not reach a steady state at times $t=11t_{r}$.
For the two-phase region, we found that the degree distribution follows a double power law distribution with two critical exponents $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$ for $k>6$. We noted that $\gamma_{1}<\gamma_{2}$, being $\gamma_{1}$ more stable over time. Moreover, we analyzed the average clustering coefficient when clustered particles are present. Because of the disk geometry, we analytically showed that this system tries to reach values around $\bar{C}=0.4$ for large packing fractions. The unexpected scaling for the average path length $l_{G}=C(\mbox{Pe})\phi^{1.2}$ is discussed, providing a simple tool to explain our results in the inner region. Our work opens exciting avenues for future research, using complex networks to characterize phase transitions in living organisms, from bacteria and ants to humans. We anticipate our study’s appeal to ecological systems, transport phenomena, multi-agent-based models, and dynamic complex networks. Furthermore, our findings introduce new possibilities for controlling and tracking clusterization in living and artificial active matter. The analytical predictions presented here can serve as a foundation for studying systems with complex interactions, including polarization, alignment, and taxis.
Acknowledgments
The authors are grateful for the helpful feedback given Pablo de Castro. I.S and F.G.-L. have received support from the ANID – Millennium Science Initiative Program – NCN19 170, Chile. F.G.-L. was supported by Fondecyt Iniciación No. 11220683. A.N. acknowledges the financial support from the project Fondecyt Iniciación No. 11220266.

### v2

@@
-Dynamical complex networks for understanding motility induced phase separation in Active Brownian Particles
+Emerging double power-law through dynamical complex networks in motility-induced phase separation in Active Brownian Particles
 Abstract
-We investigate the dynamics of active Brownian particles (ABP) through a temporal complex network framework, focusing on degree probability distribution, average path length, and clusterization across Motility-Induced Phase Separation (MIPS) regions.
-In the single MIPS phase, particle encounters resemble a random graph.
-Transitioning toward the MIPS frontier introduces a non-power-law distribution, while the phase-separated region exhibits a double power-law distribution.
-Numerical analysis indicates a shift from random to power-law complex networks based on the system’s MIPS-phase position.
-In summary, our numerical and theoretical analysis reveals the phase transition of the system between two states and how this transition manifests itself in the complex network, showing an evident change between a random network and a network that tends to have scale-free behavior.
-1 Introduction
+We investigate the behavior of active Brownian particles (ABP) within a temporal complex network framework approach.
+We focused on the node degree distribution, average path length, and average clustering coefficient across the Péclet number and packing fraction region.
+In the single phase or gas region, particle interactions mirror a random graph, and the average number of unique interactions display a non-monotonic behavior with the packing fraction.
+As we ventured toward the Motility-Phase Separation (MIPS) frontier, a hybrid distribution with a time-evolving pattern emerged, combining Gaussian and power law components.
+Moreover, we discovered a double power-law distribution in the phase-separated region, with two characteristic slopes symbolizing the emerging gas and solid regions.
+Our approach involved various numerical and theoretical analyses to capture the role of the packing fraction and Péclet number in the topological properties of the dynamical complex network that arises during the ABP dynamics.
+These findings provide a robust understanding of a phase transition between two states and how this transition manifests as a random to scale-free behavior.
+preprint: AIP/123-QED
+I Introduction
 Active Brownian Particles (ABP) serve as a valuable framework for investigating the non-equilibrium dynamics and statistical properties of self-propelled agents, whether they are living organisms or artificial entities romanczuk2012active ; bechinger2016active .
@@
 This phase separation is a fascinating emergent phenomenon where particles collapse into clusters although natural repulsion forces are present.
-The exploration of ABP not only deepens our understanding of self-organization and emergent phenomena but also opens up promising avenues for harnessing their unique properties in various applications and technological advancements gompper20202020 and building a bridge between physical models and biological systems.
+The exploration of ABP not only deepens into the understanding of self-organization and emergent phenomena but also opens up promising avenues for harnessing their unique properties in various applications and technological advancements gompper20202020 and building a bridge between physical models and biological systems.
 The complexity of ABP dynamics and its statistical properties offer novel opportunities for deeper complex network analysis to better understand the role of particle connections during the dynamics.
@@
 Our approach suggests that a random graph can characterize the dilute phase while a double power law distribution describes the MIPS region.
-2 Active Brownian Particles (ABP) and motility-phase-separation
+II Active Brownian Particles (ABP) and motility-phase-separation
 We consider $N$ self-propelled disk-like particles with position vectors $\mathbf{r}_{i}(t)=[x_{i}(t),y_{i}(t)]$ (center of the disk) and orientation vectors $\hat{\mathbf{u}}_{i}(t)=[\cos\theta_{i}(t),\sin\theta_{i}(t)]$ that obey overdamped Langevin dynamics in a flat 2D surface,
-$\dot{\mathbf{r}}_{i}$ $=$ $u_{0}\hat{\mathbf{u}}_{i}(t)+\bm{\xi}_{i,T}(t)/\gamma_{T}-(1/\gamma_{T})\nabla_{\mathbf{r}_{i}}U,$ (1)
+$\dot{\mathbf{r}}_{i}$ $=$ $u_{0}\hat{\mathbf{u}}_{i}(t)+(1/\gamma_{T})(\bm{\xi}_{i,T}(t)-\nabla_{\mathbf{r}_{i}}U),$ (1)
 $\dot{\hat{\mathbf{u}}}_{i}$ $=$ $(1/\gamma_{R})\bm{\xi}_{i,R}(t)\times\hat{\mathbf{u}}_{i}$ (2)
-where $\gamma_{T},\gamma_{R}$ are the Stokes friction coefficients associated with translation (T) and rotation (R) in the $x-y$ plane, respectively, $u_{0}$ represents the self-propelled velocity with constant magnitude see Figure 1.
+where $\gamma_{T},\gamma_{R}$ are the Stokes friction coefficients associated with translation (T) and rotation (R) in the $x-y$ plane, respectively, $u_{0}$ represents the self-propelled velocity with constant magnitude see Fig.
+1.
 Also, in Eqs.(1) and (2), $\bm{\xi}_{T}(t)$ and $\bm{\xi}_{R}(t)$ are, respectively, the translational and rotational, Gaussian, white noises with zero mean, $\langle\bm{\xi}_{T}(t)\rangle=\langle\bm{\xi}_{R}(t)\rangle=\mathbf{0}$, and time correlation $\langle\bm{\xi}_{T}(t_{1})\bm{\xi}_{T}(t_{2})\rangle=D_{T}\delta(t_{1}-t_{2})$ and $\langle\bm{\xi}_{R}(t_{1})\bm{\xi}_{R}(t_{2})\rangle=D_{R}\delta(t_{1}-t_{2})$, where we can define $D_{T}$ and $D_{R}$ as the translation and rotation diffusion parameters.
@@
 with $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ is the distance between particles, $\sigma$ the particle diameter, $C_{r}=2^{1/6}\sigma\approx 1.1225\sigma$ is defined here as the contact radius, and $\varepsilon$ is the energy interaction strength between particles.
-We set $\varepsilon=3$ during our simulations unless we stated otherwise.
-Figure 1: Time evolution for a system of active Brownian particles exhibiting Motility-induced phase separation: In the top-left, a single ABP is shown with orientation vector $\hat{u}_{i}$ and radius $R$.
-The time evolution is shown at three different times, showing a transition from a single phase in $t=0.02t_{r}$ and two coexisting phases in $t=5.85t_{r}$, at $\text{Pe}=140$ and $\phi=0.5$, with $L=100$ and $t_{r}=1/D_{R}$.
+We set $\varepsilon=3$ unless we stated otherwise during our simulations.
+Figure 1: Time evolution for a system of active Brownian particles exhibiting Motility-induced phase separation: In the top-left inset, a single ABP is shown with orientation vector $\hat{u}_{i}$ and radius $R$.
+The time evolution is shown at four different times, showing a transition from a single phase in $t=0.0t_{r}$ and two coexisting phases in $t=6.0t_{r}$, at $\text{Pe}=140$ and $\phi=0.5$, with $L=100$ and $t_{r}=1/D_{R}$.
 In this study, we conducted numerical simulations involving a collection of ABP moving within a square of area $L\times L$.
@@
 The system starts with a homogeneous particle distribution and orientations at $t=0$ exhibiting a single-phase behavior (or fluid-like state su2023motility ), which physically describes ABP randomly moving without particle agglomeration.
-However, if the packing fraction and Péclet number are larger than some critical values, the system evolves into a two-phase state represented by a solid and gas phase, as illustrated in Figure 1.
+However, if the packing fraction and Péclet number are larger than some critical values, the system evolves into a two-phase state represented by a solid and gas phase, as illustrated in Fig 1.
 In this phase, particles agglomerate, and cluster formation becomes an emergent phenomenon.
@@
 $f_{c}=\frac{4\phi\text{Pe}-3\pi^{2}\kappa}{4\phi\text{Pe}-6\sqrt{3}\pi\kappa\phi}$ (5)
+Figure 2: Analytical representation for the phase diagram in the $\text{Pe}-\phi$ plane plotted from the analytical model using Eq.
+(5), with $\kappa=4.7$ redner2013structure .
+The white dots represent the pairs $(\text{Pe},\phi)$ parameters studied across this work.
+The different colors represent the observed networks.
+White dots represent the single phase where a random graph is observed.
+The yellow dots represent the hybrid phase, and the red dots represent the two-phase region where a double power-law distribution is recovered.
 In Fig.
 2, we analytically reproduced the phase diagram in terms of the packing fraction and the Péclet number using $\kappa=4.7$ as a fit parameter redner2013structure .
-The pairs $(\text{Pe},\phi)$ studied in this work are represented by the white dots of Fig.
+The pairs $(\text{Pe},\phi)$ studied in this work are represented by dots in Fig.
 2.
-For $f_{c}=0$, we can say that the system is in the single-phase, while for $f_{c}\neq 0$, the system undergoes a phase transition to the two-phase state.
+For $f_{c}=0$, we can say that the system is in the single-phase (white dots), while for $f_{c}\neq 0$, the system undergoes a phase transition to the two-phase state (red dots).
 The system is considered steady when the phase separation shows a bimodal probability distribution for different local packing fractions redner2013structure ; stenhammar2014phase .
@@
 In this work, we aim to study the time evolution by generating a temporal averaged network during the phase transition.
-Figure 2: An analytical representation for the phase diagram in the $\text{Pe}-\phi$ plane plotted from the analytical model using Eq.
-5, with $\kappa=4.7$.
-The white dots represent the pairs $(\text{Pe},\phi)$ parameters studied across this work.
-2.1 Simulation methods and parameters
+II.1 Simulation methods and parameters
 We employed standard Brownian dynamics methods to numerically solve the governing Eqs.(1) and (2) for $N\in[955,1274,1592,1910,2229]$ particles within a simulation box with a length of $L=100$ and periodic boundary conditions.
@@
 Additionally, we adopted $t_{r}$ as the unit of time.
-The equations of motion, Eqs.(1) and (2), were integrated with a time step of $\Delta t=3\times 10^{-5}t_{r}$, and simulation times were set to $T=12t_{r}$.
+The equations of motion, Eqs.(1) and (2), were integrated with a time step of $\Delta t=3\times 10^{-5}t_{r}$, and simulation times were set to $T=30t_{r}$.
 Two dimensionless values parameterize the system: the packing fraction $\phi$ and the Péclet number, which are varied in the intervals $\phi\in[0.3-0.7]$ and $\text{Pe}\in[80,140]$, respectively (see Fig.
 2).
-We maintain a constant rotational diffusion coefficient $D_{R}=0.15$, translation diffusion coefficient $D_{T}=0.2$, and diameter $\sigma=2R$ with $R=1$.
-3 Complex Networks for ABP
+We maintain a constant rotational diffusion coefficient $D_{R}=0.15$, translation diffusion coefficient $D_{T}=0.2$, the diameter $\sigma=2R$ with $R=1$, and the Stokes translational and rotational friction coefficients are $\gamma_{T}=6\pi$ and $\gamma_{R}=8\pi$ respectively.
+III Complex Networks for ABP
 A network is defined as a group of nodes connected between them by edges.
@@
 We build a time network by following the dynamics of the particle connections in a specific time window.
+The network is built considering a fixed number of nodes $N$ and an increasing number of connections.
 To have enough statistics, we define a window with sixty data points, and we count the connections between nodes each time; if two nodes are connected only once in a one-time step or if two nodes are connected along the sixty steps, the value of the connections in the adjacency matrix is one in that time window.
@@
 3.
-Figure 3: From left to right: A scheme for particle encounters when a pair interaction occurs.
-The blurred area represents the interaction region with radius $C_{r}$.
-Particles (nodes) generate a link if their centers enter each other’s interaction region, as shown in the middle panel.
-A representation of the time evolution for the network formation, with time windows of length $W=0.45t_{r}$ and overlap $t_{0}=0.225t_{r}$ between two consecutive windows.
-3.1 Adjacency matrix
+III.1 Adjacency matrix
 In order to analyze the system, we first create the adjacency matrix $A$, which is used to represent the network.
@@
 The latter means that each network accounts for unique particle encounters within the time window.
+Figure 3: From left to right: Scheme for particle encounters when a pair interaction occurs.
+The blurred area represents the interaction region with radius $C_{r}$.
+Particles (nodes) generate a link if their centers enter each other’s interaction region, as shown in the middle panel.
+A representation of the time evolution for the network formation, with time windows of length $W=0.45t_{r}$ and overlap $t_{0}=0.225t_{r}$ between two consecutive windows.
 From the adjacency matrix, we can compute all the necessary properties of the network.
+As we know, the study of complex networks can shed light on the global behavior of a system composed of small and separated elements.
 We will focus on calculating the degree distribution, network average clustering coefficient, and average path length.
-3.2 Degree distribution
-The degree $deg(p_{i})$ of each particle $p_{i}$ is the number of connections of a node, and it can be computed by summing over the corresponding column or row of the adjacency matrix:
+These metrics give good insights into describing, characterizing, and understanding the time evolution of the structure of a complex network.
+III.2 Degree distribution
+As a first manner to characterize these complex networks, we compute the degree of the nodes.
+We define the degree $deg(p_{i})$ of each particle $p_{i}$ as the number of connections of a node, and it can be computed by summing over the corresponding column or row of the adjacency matrix:
 $deg(p_{i})=\sum_{j=1}^{N}A_{ij}=\sum_{j=1}^{N}A_{ji},$ (6)
@@
 Here $n_{k}$ is the total number of particles with degree $k$, and $N$ is the number of particles in the system.
-Depending on the topology of the complex network, the degree distribution could be binomial (random graphs), power-law (scale-free), among others.
-In this work, we shall discuss the role of the Gaussian distribution and double power law models in the context of MIPS.
-3.3 Network average clustering coefficient
-The clustering coefficient $C_{i}$ of each particle is a measure of the tendency to form clusters of this node, and it is defined by
+Depending on the topology of the complex network, the degree distribution could show different behaviors Erdos ; barabasi1999 ; albert2002statistical ; newman2003 : it could be binomial (random graphs) or power-law (scale-free).
+Real networks usually show a scale-free behavior.
+To explain this behavior of a complex network built with the World Wide Web, Réka Albert and Albert Barabási barabasi1999 ; albert2002statistical ; barabasi2011 used a model for random networks given a preferential probability of attachment for the new connections.
+In this work, we shall discuss the role of the Gaussian distribution and double power law models in the context of MIPS to understand the global structure of these complex networks.
+Also, the degree distribution has been shown to be useful to identify phase transitions in real networks pasten2016 .
+III.3 Network average clustering coefficient
+Another characteristic of real networks is the small-world phenomenon.
+Watts and Strogatz watts1998 show that random networks do not have this behavior in which two nodes of the complex network are connected between them through small steps.
+In order to give a measure of the small world behavior, we can define a global clustering coefficient (for all the nodes of the network) and a local clustering coefficient (for each node).
+The values of the clustering coefficient allow us to identify these small-world properties.
+Both random networks and ordered lattices do not show a small-world behavior.
+The clustering coefficient Newman_2000 ; newman2003 $C_{i}$ of each particle is a measure of the tendency to form clusters of this node, and it is defined by
 $C_{i}=\frac{1}{k_{i}(k_{i}-1)}\sum_{j,k}A_{ij}A_{jk}A_{ki},$ (8)
 where $A$ is the adjacency matrix, and $k_{i}$ is the degree of the particle.
-From this, we can calculate the network average clustering coefficient $\bar{C}$, which is the trend of the network to form clusters, and it is defined by
+From this, we can calculate the network average clustering coefficient watts1998 $\bar{C}$, which is the trend of the network to form clusters, and it is defined by
 $\bar{C}=\frac{1}{N}\sum_{i=1}^{N}C_{i},\quad\quad 0\leq\bar{C}\leq 1.$ (9)
-3.4 Average path length
+III.4 Average path length
+To identify the small-world phenomenon, we need an additional measure: the distance between two nodes, called average path length Newman_2000 .
+The small-world model shows an increase in the log scale of the average path length and the clustering properties at the same time.
 The average path length is the rate between the shortest number of steps between two nodes and the shortest paths between all the pairs of nodes, and it is defined by Newman_2000 ,
@@
 Here, $d(v_{i},v_{j})$ is the shortest distance between two vertices $v_{i}$ and $v_{j}$, where $v_{i,j}$ are nodes of the graph.
-4 Results and Discussion
+The small-world model shows a single length scale that depends on the probability of the connections between nodes newman_watts1999 .
+IV Results and Discussion
 We studied three different regimes of the ABP according to the phase diagram of Fig.
-2: (i) one-phase (violet), (ii) transient phase (blue-green), and (iii) two-phase (orange-red).
-As discussed in the following sections, these three regions will exhibit different topological properties, which are not reported in the literature to the best of our knowledge.
-4.1 One-phase region: Random Graph
+2: (i) one-phase (white circles), (ii) transient phase (yellow circles), and (iii) two-phase (red circles).
+As discussed in the following sections, these three regions will exhibit different topological properties.
+IV.1 One-phase region: Random Graph
 In the one-phase region (violet in Fig. 2), we found that the random motion of ABP can be characterized by a Gaussian function for the degree distribution
 $P^{\rm G}(k)={\frac{1}{\sigma_{\rm st}\sqrt{2\pi}}}\mbox{exp}\left[-\frac{1}{2}\left(\frac{k-\mu}{\sigma_{\rm st}}\right)^{2}\right],$ (11)
-where $\sigma_{\rm st}$ is the standard deviation, and $\mu$ is the average number of unique encounters of particles in a time window $W$.
-We have a dilute system for packing fraction numbers $\phi<0.4$, which allows us to analyze the average behavior of the ABP system using the kinetic theory of gases.
-As explained in Ref.
-norambuena2020understanding , in an active media with $N$ moving particles in a dilute system, the mean free path is $\lambda=\sqrt{2}u_{0}\tau_{c}$, where $\tau_{c}$ is the mean time between encounters.
-During a time interval $\tau_{c}$, we can establish the probability constraint $p_{c}=A_{\rm sw}/L^{2}\approx 1$, where $A_{\rm sw}=N(2\lambda a+\pi a^{2})$ is the total area swept area of the $N$ particles, and $L^{2}$ is the available area.
-From this argument, it follows that $\lambda=(\pi/4)\phi^{-1}(1-\phi)\sigma=\sqrt{2}u_{0}\tau_{c}$.
-Then, the average number of encounters can be estimated as $\mu=W/\tau_{c}$, since, on average (by definition), particles have one encounter in the time interval $\tau_{c}$.
-We found:
-$\mu(\phi,\text{Pe})={4\sqrt{2}\over\pi}\left({\phi\over 1-\phi}\right)WD_{R}\text{Pe},$ (12)
-Figure 4: Degree distribution $P(k)$ for systems in the one-phase region at $\text{Pe}=80$ (circles) and $\text{Pe}=140$ (diamonds) at time $t=0.45t_{r}$ (top) and at $t=11.0t_{r}$ (bottom).
+[…]
+$\mu(\phi,\text{Pe})=\Lambda(\text{Pe})\left({\phi\over 1-\phi}\right)(1-\chi\phi)P(\lambda),$ (12)
+Figure 4: (a), (b) Time evolution for the degree distribution $P(k)$ for systems in the one-phase region at $\text{Pe}=20$ (triangles) and $\text{Pe}=140$ (circles) at time $t=0.45t_{r}$ and at $t=30.0t_{r}$ for different packing fractions $\phi$.
 The distributions show a Gaussian tendency; each system exhibits distinct mean and standard deviation values.
-Note that all systems maintain a Gaussian distribution, and only the system at $\phi=0.4$ shows a change over time.
-Both insets (top and bottom) show the $\log-\log$ plot of the systems at the respective time.
-where $W=0.45t_{r}$ is fixed for $D_{R}=0.15$.
+Note that all systems maintain a Gaussian distribution.
+The inset shows the average number of unique encounters between particles $\mu$, for different Pe and $\phi$, extracted from (c).
+(c) The average number of unique encounters of particles $\mu$ for different Péclet numbers in the single-phase, dashed lines show the analytical value following Eq.
+(12).
+where $\Lambda(\text{Pe})=4\sqrt{2}WD_{R}\text{Pe}/\pi$, $P(\lambda)=1-e^{-(\lambda-\sigma)/\lambda}$, $\lambda=\lambda(\phi)$, $W=0.45t_{r}$, $D_{R}=0.15$, and we use $\chi=0.2$ to have a better fit.
 In Fig.
-4, we illustrate the evolution of the degree distribution $P(k)$ measured from Eqs.
-6 and 7 for a system at two time lapses, $t=0.45t_{r}$ and $t=11.0t_{r}$.
-Two Péclet numbers are considered: $\text{Pe}=80$ represented by circles and $\text{Pe}=140$ represented by diamonds.
-Different packing fractions are depicted with orange and blue colors, corresponding to $\phi=0.3$ and $\phi=0.4$, respectively.
-Most importantly, using our model given in Eq.
-12, we obtain $\mu(\phi=0.3,\text{Pe}=80)\approx 10.3$ and $\mu(\phi=0.3,\text{Pe}=140)\approx 16.2$, whose numerical values are close to those obtained by simulations ($10.51$ and $16.75$), see Fig.
-5.
-This simple model can give good estimations of the central value of the Gaussian distribution $P^{\rm G}(k)$ given in equation 11, which is valid for the one-phase region.
-Figure 5: Standard deviation $\sigma_{st}$ (top) and mean $\mu$ (bottom) values over time of Gaussian fits for the degree distribution of systems in the one-phase region (Fig.
-4).
-Both $\sigma_{st}$ and $\mu$ remain virtually constant over time and only the mean for the system at $\phi=0.4$ shows an appreciable decrease.
-4.2 Transient Region: Hybrid phase
-We define the transient region as the intermediate phase between the one-phase and two-phase regions (blue-green in Fig.
+4 (a), (b), we show the time evolution for the Gaussian distributions of $P(k)$ for different packing fractions and Péclet numbers.
+Using our model given in Eq.
+(12), we obtain $\mu(\phi=0.1,\text{Pe}=140)\approx 7.118$, $\mu(\phi=0.2,\text{Pe}=140)\approx 13.456$, and $\mu(\phi=0.3,\text{Pe}=140)\approx 16.686$ whose values are close to those obtained by our simulations ($6.804$, $12.659$, $16.773$), see Fig.
+4 (c).
+This simple model can give good estimations of the central value of the Gaussian distribution $P^{\rm G}(k)$ given in Eq.
+(11) in the regime $\phi<\phi_{c}=0.41$.
+The inset in Fig.
+4 (b) resumes the obtained values for $\mu$ in our simulations.
+Showing a non-monotonic behavior, where $\mu$ increases with $\phi$ for dilute concentrations and $\mu$ decreases for denser systems.
+IV.2 Transient Region: Hybrid phase
+We define the transient region as the intermediate phase between the one-phase and two-phase regions (blue in Fig.
 2).
@@
 In Fig.
-6, we observe the degree distribution, average path length ($l_{G}$), and average clustering coefficient ($\bar{C}$) at times $t=0.45t_{r}$ and $t=11t_{r}$.
+5, we observe the time evolution for the degree distribution at times $t=0.45t_{r}$ and $t=30t_{r}$.
 Interestingly, the initially random graph (Gaussian distribution) evolves to a hybrid distribution with two components (Gaussian and power law).
 This hybrid behavior tells us that the system is trying to evolve from a random to a scale-free graph, but the values for the packing fraction and Péclet are not large enough.
-Moreover, at $t=11t_{r}$, the system does not reach the steady state, which is not the case for the one-phase and two-phase regions.
+Moreover, at $t=30t_{r}$, the system does not reach the steady state, which is not the case for the one-phase and two-phase regions.
 The latter tells us that this hybrid phase is more complex to analyze regarding its geometrical and topological properties since the system remains out of equilibrium.
-Figure 6: Top: Degree distribution for the case at $\text{Pe},\phi=80,0.4$ at the initial time $t=0.45t_{r}$ (right) and time $t=11.0t_{r}$ (left).
-The inset shows the same data plotted in $\log{}-\log{}$.
-Bottom: Average path length (left ) and average clustering coefficient (left) over time of the system.
-Both show an increasing behavior over time until $\approx 8t_{r}$.
-4.3 Two-phase region: Phase separation
+In Fig.
+2, we colored the data points where we observed this regime in yellow circles.
+Figure 5: Time evolution for the degree distribution at the hybrid phase.
+At the initial time $t=0.45t_{r}$, a Gaussian distribution is observed for the different systems; meanwhile, a hybrid distribution is observed at time $t=30.0t_{r}$.
+The inset shows the same data plotted in $\log{}-\log{}$ scale.
+IV.3 Two-phase region: Double power law distribution
 At more significant packing fractions and Péclet numbers, i.e., $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$, the system enters the inner region of Fig.
 2, where MIPS is manifested.
-From the point of view of the ABP, as time evolves, the initially random distribution (liquid-like state) evolves to a state where a gas-like and a solid-like phase cohabit in the system (green-red regions in Fig.
+From the point of view of the ABP, as time evolves, the initially random distribution (liquid-like state) evolves to a state where a gas-like and a solid-like phase cohabit in the system (green-yellow regions in Fig.
 2).
 In terms of the complex network, the degree distribution of the system evolves from a Gaussian distribution at short times ($t=0.45t_{r}$) towards a double power law distribution for larger times ($t\geq 11t_{r}$); see Fig.
-7(a)(b).
+6 (a),(b).
 Notably, when the degree distribution is plotted in the log-log scale, two slopes adequately describe the distribution $P(k)$, see Fig.
-7(c)(d).
+6 (c), (d).
 This double power law distribution has the form
@@
 The value $K=6$ is connected with the 2D disk packing problem, where each disk has exactly six nearest neighbors in an ideal cluster and corresponds to the peak of $P(k)$, see Fig.
-7(b).
+6 (b).
 Moreover, the critical values for $k_{c}$ depend on the clustering coefficient and Péclet number.
 The first slope $\gamma_{1}$ (red dashed line of Fig.
-7) changes from values close to $1.5$ at $\phi=0.5$ towards values close to $3.7$ at $\phi=0.7$.
+6 (c), (d)) changes from values close to $1.5$ at $\phi=0.5$ towards values close to $3.7$ at $\phi=0.7$.
 The second slope $\gamma_{2}$ (black dashed lines) varies from values close to $12$ to $17$.
@@
 To better understand this, we plot the time evolution in Fig.
-8, where it is clear that the slope value and, therefore, the cluster solid-like region is determined strongly by the system packing fraction $\phi$.
+7, where it is clear that the slope value and, therefore, the cluster solid-like region is determined strongly by the system packing fraction $\phi$.
 On the contrary, the $\gamma_{2}$ exponent is related to less probable encounters related to the gas region, and we will not delve into that aspect.
@@
 We note that $\gamma_{1}$ is more stable over time, and the packing fraction dominates its steady value in some inverse relation.
-On the contrary, $\gamma_{2}$ shows more oscillations, which can be explained because the clustering formation is unstable for the values of $(\phi,\mbox{Pe})$.
+On the contrary, $\gamma_{2}$ shows more fluctuations, which can be explained because the clustering formation is unstable for the values of $(\phi,\mbox{Pe})$.
 Table 1: Comparison between critical exponents for the double power law distribution in different models.
@@
 (a) $\log{}-\log{}$ plot shows a clear Gaussian distributions at time $t=0.45t_{r}$.
-(b) degree distribution at $t=11.0t_{r}$ shows an evolution in the distribution where the degree $k=6$ for each curve is the one with the highest probability.
-Figures (c) and (d) show two distributions of the figure (b) zoomed in at $\text{Pe},\phi=(80,0.7)$ and $\text{Pe},\phi=(140,0.5)$ respectively, where the red and black dashed lines are the best linear fit for the first and second power law decay respectively for each distribution.
-Figure 8: Slopes of the power law decay in the degree distribution for phase-separated systems.
-The initial data point occurs when the degree $k=6$ has the highest probability, indicating distinct curves for each system.
-Consequently, we observe different transition times.
-Notably, around $t\approx 8t_{r}$, the value of $\gamma_{1}$ (top) stabilizes consistently across all systems.
-In contrast, the values for the second slope, $\gamma_{2}$ (bottom), appear noisy and do not seem to stabilize.
-4.3.1 Role of the average clustering coefficient
-We can characterize this phase separation using the network average clustering coefficient $\bar{C}$ given in Eq.
+(b) degree distribution at $t=12.0t_{r}$ shows an evolution in the distribution where the degree $k=6$ for each curve is the one with the highest probability.
+(c) and (d) shows two distributions of figure (B) zoomed in at $\text{Pe},\phi=(80,0.7)$ and $\text{Pe},\phi=(140,0.5)$ respectively, where the red and black dashed lines are the best linear fit for the first and second power law decay respectively for each distribution.
+Figure 7: Time evolution for the slopes fitted for the power law distribution (13), in the phase-separated systems.
+IV.3.1 Role of the average clustering coefficient
+According to our simulations, particles agglomerate when $(\phi,\mbox{Pe})$ cross some critical region, and a natural tendency to form clusters appears.
+In the language of complex networks, we can elaborate on the connection between this phase separation using the tendency to form clusters from a topological point of view.
+We encoded the topological information into a dynamical adjacency matrix constructed upon unique encounters in a time interval.
+Then, we can use the average clustering coefficient presented in Eq.
+(9) to analyze the global property of such connections to learn how to interpret the inherent hexagonal packing of the solid disks used in our ABP model.
+In the upper panel of Fig.
+8, we show the temporal behavior of the average clustering coefficient $\bar{C}$ varying the packing fraction ($\phi$) and Péclet number (Pe) and covering different points of the phase diagram illustrated in Fig.
+2.
+In all cases, the natural tendency for a fixed value of Pe is to have higher values of $\bar{C}$ as the packing fraction increases, where $0.1\leq\bar{C}\leq 5.4$.
+Also, the persistent motion represented by the Péclet number shows that for $\phi\geq 0.5$, an increment of Pe increases does not affect $\bar{C}$ in the long-time behavior, see upper right panel of Fig.
+8.
+To have an intuition about the higher values of $\bar{C}$, we show in Fig.
+9 the average clustering coefficient when particles agglomerate, forming perfect hexagonal clusters.
+In this ideal cluster, the asymptotic behavior is to reach a value $\bar{C}=0.4$, which can be understood as follows.
+Let be $\bar{C}_{hex}(N)$ the average clustering coefficient for a hexagonal arrangement with $N$ particles.
+It is not difficult to show that for $N$ particles, the fractions of particles inside the borders are given by $f_{i}=(1+3N_{r}(N_{r}-1))/(1+3N_{r}(N_{r}+1))$, where $N_{r}=[(1+4/3(N-1))^{1/2}-1]/2$.
+Moreover, each particle inside the cluster has a clustering coefficient $C_{i}=6/15$.
+For the particles at the border, we have particles in the corners with clustering coefficient $C_{1}=2/3$, and for the rest of the particles, we have $C_{2}=1/2$.
+The fraction of particles at the corners is $f_{c}=1/N_{r}$.
+Based on this, we found the analytical expression:
+$\bar{C}_{hex}(N)=\left[f_{c}(N)C_{1}+(1-f_{c}(N))C_{2}\right](1-f_{i}(N))+C_{i}f_{i}(N),$ (14)
+which gives us the formula to calculate the average cluster coefficient for an ideal hexagonal cluster.
+This expression perfectly matches the numerical data points presented in Fig.
 9.
-In Fig.
-9, we plot $\bar{C}$ as a function of time for different values of the packing fraction and Péclet number for the frontier and inner regions of the phase diagram illustrated in Fig.
-2.
-To better understand the observed values of $\bar{C}$, we can elaborate a simple theoretical analysis.
-The clustering coefficient can be computed as $C=\lambda_{G}(v)/\tau_{G}(v)$, where $\lambda_{G}(v)$ is the number of triangles with one vertex on $v$ and $\tau_{G}(v)$ is the number of triples with one vertex on $v$ ($v$ is a vertex of the graph $G$).
-For a simple hexagonal cluster composed of seven particles (see inset of Fig.
-10 (left)), we can calculate the clustering coefficient for one particle inside the cluster ($C_{i}$) and one particle in the border ($C_{b}$).
-Using this reasoning, we obtain $C_{i}=6/15=0.4$ and $C_{b}=2/3\approx 0.6667$.
-Let be $\bar{C}_{hex,N_{r}}$ the average clustering coefficient for a hexagonal arrangement with $N_{r}$ concentric hexagonal rings.
-For $N_{r}$ hexagonal rings, we have $6N_{r}$ particles in the border and $1+3N_{r}(N_{r}+1)$ particles in total.
-Therefore, the fraction of particles in the border and inside the cluster are $f_{b}(N_{r})=6N_{r}/(1+3N_{r}(N_{r}+1))$ and $f_{i}=(1+3N_{r}(N_{r}-1))/(1+3N_{r}(N_{r}+1))$, respectively.
-Thus, we can define the average clustering coefficient as,
-$\bar{C}_{hex,N_{r}}=f_{b}(N_{r})C_{b}+f_{i}(N_{r})C_{i}={2\over 15}\left[5-2f_{i}(N_{r})\right],$ (14)
-where we have used $f_{b}+f_{i}=1$.
-For instance, for $N_{r}=1$ (seven particles) we get $\bar{C}_{hex,7}=(1/7)C_{m}+(6/7)C_{b}\approx 0.6286$ (first value of Fig.
-10 (left)).
-However, as one increases the fractions of particles inside the cluster ($f_{i}\rightarrow 1$), we get $\bar{C}_{hex,N_{r}}\rightarrow C_{i}=0.4$ (asymptotic value of Fig.
-10 (left)).
-Also, it is important to mention that three disks forming a minimal cluster have an average cluster coefficient $\bar{C}=1$ (maximum value).
-In Fig.
-9, we note that in the frontier region of the MIPS phase diagram ($\phi=\{0.3,0.4\}$ and Pe =$=\{80,140\}$), where small triangular clusters randomly appear, we observe $0.2\lesssim\bar{C}\lesssim 0.35$, and we associate these values for very low probability events (probabilities $\sim(0.2-0.35)$) of forming small triangular clusters.
-However, as we enter the inner region of the MIPS phase diagram ($\phi=\{0.5,0.6,0.7\}$ and Pe =$=\{80.140\}$), finite hexagonal clusters are dominant, explaining the observed values $0.35\lesssim\bar{C}\lesssim 0.42$ at large times oscillates around $0.4$.
-A steady state value $\bar{C}\approx 0.35$ can be explained by the fact that a perfect hexagonal cluster with a clustering coefficient of 0.4 can be penalized by the fraction of particles with respect to the total number of particles.
-A value $\bar{C}\approx 0.42$ can be understood as the reference value 0.4 with a small contribution of small triangular clusters.
-Figure 9: Top: Network average clustering coefficient for systems in the frontier region (top) and the inner region (bottom).
-The values for $\bar{C}$ for one-phased systems(top) at $\phi=0.3$ have virtually no change over time and slowly increase at $\phi=0.4$.
-In contrast, $\bar{C}$ for the two-phased systems increases rapidly at first and then stabilizes around $t=8t_{r}$, showing a tendency to $\bar{C}=0.4$.
-Figure 10: Left: Network average clustering coefficient of identical circles in a hexagonal packing arrangement, using $C_{r}=2^{1/6}\sigma$ ( black circles).
+Most importantly, this tells us that as $N$ increases, the system stabilized around $\bar{C}=0.4$ since $\lim_{N\rightarrow\infty}\bar{C}_{hex}(N)=6/15=0.4$.
+IV.3.2 Role of the average path length
+The average path length defined in Eq.
+(10) was measured through all regions, and their values are shown in the lower panel in Fig.
+8.
+We observe the time evolution for the network average path length in the case of a system with a single phase characterized by a Gaussian degree distribution $l_{G}$ has a constant value for different Pe numbers and different packing fractions $\phi$.
+Meanwhile, on the transient and the two-phase region, the $l_{G}$ value increases slowly in time, reaching a constant value when $t/t_{r}\geq 10$ for the hybrid phase and in $t/t_{r}\geq 8$ in the two-phase region, see Fig 8 (e), (f) respectively.
+This time scale is similar to the cluster growth characteristic time reported by Stenhammar et al.stenhammar2014phase .
+To make progress into understanding this behavior in Fig.
+9, we show the average path length in the ideal case of perfect hexagonal clusters with $N$ particles, where we note that a power-law of the form $l_{G}=30\phi^{0.5}$ matches our numerical calculations shown in a magenta dashed line Fig 9.
+This relation means that while the cluster grows, the network is more connected; therefore, there are more paths connecting particles.
+Figure 8: Top: Network average clustering coefficient of identical circles in a hexagonal packing arrangement, using $C_{r}=2^{1/6}\sigma$ ( black circles).
 $\bar{C}$ approaches asymptotically to $0.4$ (dashed red line), which is the value of the clustering coefficient of the center particle in the seven-particle arrangement (inset).
-Right: Average path length for the same hexagonal packing arrangement as a function of the packing fraction $\phi$ (black circles).
+Bottom: Average path length for the same hexagonal packing arrangement as a function of the packing fraction $\phi$ (black circles).
 We see that $l_{G}$ grows as $30\sqrt{\phi}$ (red dashed line) and could represent an upper bound with respect to $\phi$ for clustered ABP systems.
-4.3.2 Role of the average path length
-The average path length defined in Eq.
-10 was measured for both regions, the frontier and inner region.
-In Fig.
-11, we observe the time evolution for the network average path length in the case of a system with a single phase characterized by a Gaussian degree distribution $l_{G}$ has a constant value for different Pe numbers, meanwhile on the transient region for $\phi=0.4$ and Pe=$80$ the value increases slowly not reaching a constant value see blue dots in Figure 11.
-In Fig.
-10, we can observe a particular scaling for the average path length.
-In the ideal case (perfect hexagonal clusters with $N$ particles), we note that $l_{G}=30\phi^{0.5}$ matches our numerical calculations shown in Figure 10.
-However, our simulations shown in Fig.
-11 suggest a different scaling for the steady state value of $l_{G}$, where $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$.
-We propose the following phenomenological expression
-$l_{G}(\mbox{Pe},\phi)=\zeta(\mbox{Pe})\times\phi^{1.3},$ (15)
-where $\zeta(\mbox{Pe}=80)=21$ and $\zeta(\mbox{Pe}=140)=8.5$ are values obtained to fit our data.
-Using this simple model, we found the following estimations $l_{G}(\mbox{Pe}=80,\phi=0.5)\approx 8.53$, $l_{G}(\mbox{Pe}=140,\phi=0.5)\approx 7.51$, $l_{G}(\mbox{Pe}=80,\phi=0.6)\approx 10.81$, $l_{G}(\mbox{Pe}=140,\phi=0.5)\approx 9.53$, $l_{G}(\mbox{Pe}=80,\phi=0.7)\approx 13.21$, $l_{G}(\mbox{Pe}=140,\phi=0.7)\approx 11.64$, which are in very good agreement with our numerical calculations.
-For the inner region, we observe a time evolution in the $l_{G}$ value where the average path length increases with particle concentration $\phi$ and reaches a constant value after $t\geq 8t_{r}$ which corresponds to the cluster growth characteristic time reported by Stenhammar et al.stenhammar2014phase .
-Figure 11: Average path length for systems in the frontier region (top) and in the inner region (bottom).
-The values for $l_{G}$ for single-phase systems(top) at $\phi=0.3$ have virtually no change over time and at $\phi=0.4$ increases slowly.
-In contrast, $l_{G}$ for the separated phase increases rapidly at first and then stabilizes around $t=8t_{r}$.
-5 Conclusions
-In this work, we have studied the dynamics of ABP using complex networks.
-We build a time-averaged network and concentrate on the evolution of the topology of the complex network across the MIPS region.
-We follow the time evolution of the degree distribution, the clustering coefficient, and the average path length, finding different structures of the complex network along the time evolution of the ABP.
-In the one-phase region, we developed a numerical and theoretical analysis to explain the Gaussian distribution for random graphs.
-Using a mean free path analysis, we found an analytical expression for the average number of unique encounters of particles in a time window $W$, showing the role of the packing fraction and Péclet number.
-We reported a transient region, where the degree distribution is hybrid in the sense that it is represented by two components, Gaussian and power laws, namely.
-Also, in the hybrid region, the average path length and average clustering coefficient do not reach a steady state at times $t=11t_{r}$.
-For the two-phase region, we found that the degree distribution follows a double power law distribution with two critical exponents $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$ for $k>6$.
-We noted that $\gamma_{1}<\gamma_{2}$, being $\gamma_{1}$ more stable over time.
-Moreover, we analyzed the average clustering coefficient when clustered particles are present.
-Because of the disk geometry, we analytically showed that this system tries to reach values around $\bar{C}=0.4$ for large packing fractions.
-The unexpected scaling for the average path length $l_{G}=C(\mbox{Pe})\phi^{1.2}$ is discussed, providing a simple tool to explain our results in the inner region.
-Our work opens exciting avenues for future research, using complex networks to characterize phase transitions in living organisms, from bacteria and ants to humans.
-We anticipate our study’s appeal to ecological systems, transport phenomena, multi-agent-based models, and dynamic complex networks.
-Furthermore, our findings introduce new possibilities for controlling and tracking clusterization in living and artificial active matter.
-The analytical predictions presented here can serve as a foundation for studying systems with complex interactions, including polarization, alignment, and taxis.
+We propose the following phenomenological expression for the asymptotic values of the average path length in terms of $(\phi,\mbox{Pe})$ plotted in Fig 9,
+$l_{G}(\mbox{Pe},\phi)=\alpha+\beta\phi^{\zeta},$ (15)
+where $\alpha$, $\beta$ and $\zeta$ are fitting parameters.
+We present the best-fit results in dashed lines of different colors in Fig.
+9 where we observe that for packing fractions $\phi<0.4$, the average path length is approximately constant at different Péclet numbers; this is represented in the fit parameter $\alpha$ which is inversely proportional to Pe.
+For denser systems $\phi>0.4$, $l_{G}$ increases following a power-law with exponents between $\zeta\approx 2-4$, meanwhile $\beta\approx 20$ for all Pe.
+This behavior with increasing $\phi$ is a fingerprint of the cluster formation.
+When a solid phase appears in the steady state of the system, $l_{G}^{\text{solid}}$ contributes significantly to the network connection, suffering a drastic behavior when the phase separation emerges at large Pe.
+Figure 9: Top: Network average clustering coefficient $\bar{C}$ as a function of the particle number $N$.
+Following a crystal shape, $\bar{C}$ reaches an analytical saturation value $\bar{C}\sim 0.4$ as the cluster grows.
+Bottom: The asymptotic value of the network’s average path length $l_{G}$ as a function of the packing fraction $\phi$.
+The star symbols represent the simulation’s results for the case of a cluster growing in terms of $\phi$.
+The dashed magenta line represents the best fit, which follows a power-law behavior.
+In symbols, we present the asymptotic $l_{G}$ values measured from Fig.
+8 in terms of $\phi$ for different Pe.
+The dashed lines represent the best fits for each case using equation (10).
+V Conclusions
+In this work, we focused on studying the behavior of Active Brownian Particles (ABP) using complex networks.
+We examined the changes in the network structure as the packing fraction and Péclet number varied, particularly in the Motility-Induced Phase Separation (MIPS) region.
+We tracked the evolution of parameters such as the degree distribution, clustering coefficient, and average path length, revealing different network structures as the ABP behavior evolved.
+In the single-phase region, we conducted both numerical and theoretical analyses to explain the Gaussian distribution for random graphs.
+By analyzing the mean free path, we derived an expression for the average number of unique particle encounters within a given time window, highlighting the influence of packing fraction and Péclet number.
+Surprisingly, even in the single-phase region, we observed a critical packing fraction ($\phi\geq\phi_{c}=\pi/(\pi+2^{13/6})\approx 0.41$) that led to a significant change in network topology.
+Above this threshold, the mean free path became smaller than the contact radius ($C_{r}$) of the Weeks-Chandler-Andersen potential, indicating the pivotal role of packing fraction in network analysis.
+We also identified a transient region where the degree distribution evolved over time, resulting in a hybrid distribution combining Gaussian and power law components.
+This region corresponded to the formation of small clusters without distinct separation.
+Our observations did not align with the analytical expression by Redner et al.
+redner2013structure but matched their numerical predictions and the critical cluster region reported by McDermott et al.
+mcdermott2023characterizing using machine learning analysis in MIPS.
+Additionally, the average path length and clustering coefficient stabilized in the hybrid region after $t\geq 11t_{r}$.
+In the two-phase region, we found that the degree distribution followed a double power law distribution with two critical exponents, $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$, for $k>6$, representing solid and gas-like phases, respectively.
+We observed that $\gamma_{1}<\gamma_{2}$, with $\gamma_{1}$ reaching a stable state over time while $\gamma_{2}$ fluctuated around an average value.
+Furthermore, we analyzed the average clustering coefficient in the presence of clustered particles.
+Due to the disk geometry of each ABP, we demonstrated that the system tended to approach values around $\bar{C}=0.4$ for large packing fractions.
+We discussed the behavior of the average path length ($l_{G}$), which increased with the packing fraction and decreased with the Péclet number and provided a heuristic model to describe its behavior across different regions.
+Our research paves the way for future studies using complex networks to characterize phase transitions in various living organisms, from bacteria and ants to humans.
+We believe our findings will be relevant to ecological systems, transport phenomena, multi-agent-based models, and dynamic complex networks.
+Additionally, our work introduces possibilities for controlling and tracking clusterization in living and artificial active matter and their impact on disease transmission.
+The analytical predictions presented in this study can form the basis for investigating systems with complex interactions, including polarization, alignment, and taxis.
 Acknowledgments
-The authors are grateful for the helpful feedback given Pablo de Castro.
+The authors are grateful for the helpful feedback given by Pablo de Castro.
 I.S and F.G.-L.